OpenStudy (anonymous):

Calculus - Determine the average value of f(x) over the interval form from x=a and x=b where f(x) = 1/x a= 1/7 and b= 7

OpenStudy (anonymous):

\[\frac{ 1 }{(b-a)}\int\limits_{a}^{b} f(x) dx\]

OpenStudy (anonymous):

I got -100/7

OpenStudy (anonymous):

result is 4.0390

OpenStudy (anonymous):

\[\frac{ 7 }{ 48 } \int\limits_{1/7}^{7} \frac{ 1 }{ .5(7)^2}- \frac{ 1 }{ .5(1/7)^2 }\]

OpenStudy (anonymous):

function is 1/x so it means to be solved in boundaries a and b

OpenStudy (anonymous):

solutions of integral is 3.891

OpenStudy (anonymous):

the rest is simple math

OpenStudy (anonymous):

can you do the rest?

OpenStudy (anonymous):

I got ln(7) - ln (1/7) which is 3.891...

OpenStudy (anonymous):

then you multiply that by \[\frac{ 1 }{ b-a }\]

OpenStudy (anonymous):

7-1/7= 6.857... which 1/8.857 * 3.891 is .5675...

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

1/6.857 not 1/8.857

OpenStudy (anonymous):

so as an exact answer it would be \[7\log 7/ 24\]

OpenStudy (anonymous):

yep typo.